The length of a rectangle is more than twice its width, and the area of the rectangle is . Find the dimensions of the rectangle.
Answer by amoresroy(361)

The dimensions of the rectangle can be computed if area of rectangle is given.
Let x = width of rectangle
2x = length of rectangle
A = area of rectangle
A = x * 2x
= 2x ^2

2x^2 = A
x^2 = A/2
x = A^(1/2) / 2^(1/2)
2x = 2A^(1/2) / 2^(1/2)